Calculate the work required to be done to stop a bus of mass 2000 kg moving at a velocity of 36 km per hour

Grade:

1 Answers

Arun
6 years ago
Dear student
 

m= 2000kg

u=36km/h = 36×1000m/3600s = 10m/s

work required to be done to stop the car 

W = change in K.E. of the car

=1/2mv² - 1/2m (0)² = 1/2mu²

putting d value v get  

W= 1/2×2000× [10]² = 100000 J 

 

Regards

Arun(askIITians forum expert)